OpenStudy (anonymous):

Find the integral xtan(4x) dx

OpenStudy (across):

Try integration by parts.

OpenStudy (anonymous):

I did and I got stuck

OpenStudy (across):

That is,\[\int u(x)v'(x)\,dx=u(x)v(x)-\int u'(x)v(x)\,dx.\]

OpenStudy (anonymous):

i already know that i just am stuck on a certain part

OpenStudy (across):

On which part are you stuck?

OpenStudy (anonymous):

I get x ln(sec(4x)/4) then im stuck after that

OpenStudy (anonymous):

i need the integral at ln(sec(4x)/4)

OpenStudy (across):

I believe you mean \(\frac14\ln\cos4x\).

OpenStudy (anonymous):

no i mean what i typed